find the general solution of the homogeneous linear equations





1.(2D²-5D-3)y=0





2. (9D²-6D+1)y=0





3. (D⁴-2D²)y=0


4. (D⁵+8D³+16D)y=0


5. (D³+8)y=0


1
Expert's answer
2021-12-07T12:22:02-0500

1.

"2k^2-5k-3=0"

"k=\frac{5\pm \sqrt{25+24}}{4}"

"k_1=-0.5,k_2=3"

"y=c_1e^{-0.5x}+c_2e^{3x}"


2.

"9k^2-6k+1=0"

"k=\frac{6\pm \sqrt{36+36}}{18}"

"k_{1,2}=1\/3"

"y=c_1e^{x\/3}+c_2xe^{x\/3}"


3.

"x^4-2x^2=0"

"x_{1,2}=0"

"x_{3,4}=\pm \sqrt 2"

"y=c_1+c_2x+c_3e^{\sqrt2x}+c_4e^{-\sqrt2x}"


4.

"x^5+8x^3+16x=0"

"x(x^4+8x^2+16)=0"

"x^2=\frac{-8\pm \sqrt{64-64}}{2}=-4"

"x_1=0,x_{2,3}=2i,x_{4,5}=-2i"

"y=c_1+c_2cos2x+c_3sin2x+x(c_4cos2x+c_5sin2x)"


5.

"x^3+8=0"

"(x+2)(x^2-2x+4)=0"

"x_{1}=-2"


"x^2-2x+4=0"

"x_{2,3}=\frac{2\pm \sqrt{4-16}}{2}=1\pm i\sqrt 3"


"y=c_1e^{-2x}+e^x(c_2cos(x\sqrt 3)+c_3sin(x\sqrt 3))"